OpenStudy (anonymous):

[4/5](to the second power) + 3 4/5 - 1 3/5 =?

OpenStudy (lgbasallote):

\[\huge (\frac 45)^2 + 3\frac 45 - 1 \frac 35\] \[\huge \implies (\frac 45 \times \frac 45) + 3 \frac 45 - 1\frac 35\] does that help?

OpenStudy (waleed_imtiaz):